# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v2

@@ 6.9 SincKAN
Figures 14 and 15 show the building blocks and a synthesized activation using a fixed truncation degree $(2N+1=11)$ and spacing $h=1.0$ over $x\in[-10,10]$. Figure 14 displays the truncated Sinc dictionary $\{\mathrm{Sinc}(\tfrac{\pi}{h}(x-kh))\}_{k=-N}^{N}$, whose elements are globally supported and oscillatory with decaying sidelobes. These atoms are linearly combined to form the learned univariate map $\varphi(x)$.
@@ 7.1 Physics Constraints & Loss Design
  • Structural & Conservation Laws. Enforce hard invariants such as Hamiltonian, quantum, or algebraic constraints. Includes SPEL-KAN [241], KAN-ETS [239], and DAE-KAN [254].
@@ 10.1 Theoretical Approximation Rates and Empirical Scaling Laws
Sobolev and Besov rates. Wang [53] proved that for $f\in W^{s}(\Omega)$, a KAN with depth $L$ and spline order $k$ satisfies
@@
corresponding to a parameter-rate $\mathcal{O}(P^{-2s/d})$, doubling the classical $\mathcal{O}(P^{-s/d})$ rate of standard neural and kernel approximators [226, 227]. Kratsios et al. [271] extend this to Besov spaces [272], showing that spline-based Res-KANs achieve the optimal Besov approximation rate on Lipschitz and fractal sets with dimension-free sample complexity. Complementing these approximation-centric results, Liu et al. [140] establish the first minimax statistical rates for KAN regression: spline-based additive and hybrid KANs achieve the optimal
$O\!\left(n^{-2r/(2r+1)}\right)$
@@ 11 Practical “Choose–Your–KAN” Guide
KANs offer a modular design space. The correct configuration depends on (i) the structure of the target function (smooth, oscillatory, discontinuous, periodic, multi-scale), (ii) computational constraints (time, memory, precision), and (iii) desired properties (accuracy, interpretability). Below is a clear, seven-step recipe synthesizing Sections 610.
@@ 12 Current Gaps and Path Forward
Despite rapid growth in publications and open-source implementations (Tables 2, LABEL:githubs), the KAN ecosystem remains fragmented. Most studies show strong empirical potential in regression, operator learning, and PDE solving, yet comparisons are inconsistent and rarely grounded in theory. The field now requires principled foundations rather than isolated “KAN vs. MLP” or “basis-vs.-basis” experiments.
@@ 13 Conclusion
Kolmogorov–Arnold Networks shift the locus of learning from linear weights to univariate basis functions placed on edges, offering a flexible alternative to conventional MLPs. This review has consolidated the rapidly expanding KAN literature, with emphasis on applications in function approximation, regression, and PDE solving. Our central observation is that the key question is not whether KANs outperform MLPs, but rather which KAN is appropriate for a given task.
A clear pattern emerges across studies: well-designed KAN variants often exceed vanilla MLP and PINN baselines in accuracy and convergence, albeit with higher per-epoch computational cost. Unlocking this performance hinges on navigating KANs’ large design space. We organized this space through the lens of basis families—local B-splines and Gaussians, global Chebyshev and Fourier series—and highlighted how normalization, grid layout, and initialization shape stability and expressivity.
We also outlined the ecosystem of techniques for enhancing KAN performance. Accuracy gains arise from physics-informed losses, adaptive sampling, domain decomposition, and hybrid sequence/attention modules. Efficiency gains are enabled by GPU-aware engineering, matrix-level optimizations, and lighter bases such as ReLU-powers. These developments, combined with emerging theory on approximation and convergence, signal a transition from exploratory experimentation to principled model design.
The practical “Choose–Your–KAN” guide in Section 11 distills these insights into a decision framework linking problem characteristics to architectural and training choices. Ultimately, the strength of the KAN framework lies in its modularity: basis families, normalization schemes, and optimization strategies can be selected and composed to encode problem-specific inductive biases. Continued progress will require deeper basis-specific analysis and standardized benchmarks, but KANs have clearly established themselves as a versatile and promising direction in neural network architecture.

### v3

@@ 5.9 SincKAN
Figures 15 and 16 show the building blocks and a synthesized activation using a fixed truncation degree $(2N+1=11)$ and spacing $h=1.0$ over $x\in[-10,10]$. Figure 15 displays the truncated Sinc dictionary $\{\mathrm{Sinc}(\tfrac{\pi}{h}(x-kh))\}_{k=-N}^{N}$, whose elements are globally supported and oscillatory with decaying sidelobes. These atoms are linearly combined to form the learned univariate map $\varphi(x)$.
@@ 6.1 Physics Constraints & Loss Design
  • Structural & Conservation Laws. Enforce hard invariants such as Hamiltonian, quantum, algebraic, or geometric symmetries. Includes SPEL-KAN [249], KAN-ETS [247], DAE-KAN [262], and the symmetry-preserving GKSN architecture of Alesiani et al. [213].
@@ 6.8 Optimization & Adaptive Training
Variational basis-size learning. Alesiani et al. [212] propose INFINITYKAN, which treats the number of basis functions in each univariate map as a latent variable optimized via mean-field variational inference. Rather than fixing basis width, Poisson-distributed variables $\lambda_{\ell}$ determine layer-wise cardinality, allowing the representational capacity to expand or contract during training. Experiments show basis-size selection is a task-dependent and learnable component of KAN optimization.
@@ 9.1 Theoretical Approximation Rates and Empirical Scaling Laws
Sobolev and Besov rates. Wang [55] showed that deep spline-based KANs achieve a Sobolev approximation rate that doubles the classical parameter-rate $\mathcal{O}(P^{-s/d})$ of standard neural networks and kernel approximators, yielding $\mathcal{O}(P^{-2s/d})$ convergence [234, 235]. This accelerated rate reflects the interaction between the compositional structure of KANs and the expressive power of high-order spline bases.
Formally, the result is obtained by combining the exact reparametrization of $\mathrm{ReLU}^{k}$ networks by KANs (Theorem 5) with sharp approximation bounds for very deep piecewise-polynomial MLPs on Sobolev spaces [55]. Through this equivalence, established approximation theory for deep $\mathrm{ReLU}^{k}$ networks transfers directly to KAN architectures, leading to the following corollary.
Corollary 1 (Sobolev approximation rate for deep KANs [55])
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with smooth boundary and $s>0$. There exists a fixed width $W_{0}=W_{0}(d)$ such that for any $f\in W^{s}(\Omega)$ and any depth $L\geq 1$, there exists a KAN $g$ with width $W_{0}$, depth $L$, grid size $G=2$, and $k$-th order B-spline basis functions satisfying
@@
where $C$ is a constant independent of $L$.
Kratsios et al. [279] extend this result to Besov spaces ([280]), showing that spline-based Res-KANs achieve the optimal Besov approximation rate on Lipschitz and fractal sets with dimension-free sample complexity. Complementing these approximation-centric results, Liu et al. [146] establish the first minimax statistical rates for KAN regression: spline-based additive and hybrid KANs achieve the optimal
$\mathcal{O}\!\left(n^{-2r/(2r+1)}\right)$
@@ 10 Practical “Choose–Your–KAN” Guide
When to prefer KANs over MLPs. KANs are most advantageous when the learning problem benefits from explicit control over the underlying function space. In practice, KANs are preferable to standard MLPs when (i) data are sparse or expensive and fast convergence is required, (ii) a shallow architecture is desired for interpretability or stability, or (iii) prior knowledge about the solution structure is available. Examples include globally smooth or oscillatory functions (Chebyshev/Fourier bases), localized or multi-scale features (splines or wavelets), and nonsmooth or discontinuous behavior (Sinc- or rational-based KANs). In contrast, when large amounts of data are available and no structural prior is known, generic MLPs may remain competitive due to their simplicity and hardware efficiency.
The correct configuration of KANs depends on (i) the structure of the target function (smooth, oscillatory, discontinuous, periodic, multi-scale), (ii) computational constraints (time, memory, precision), and (iii) desired properties (accuracy, interpretability). A clear seven-step recipe addressing these aspects is given below and is also presented in algorithmic form as Algorithm 1, synthesizing Sections 59.
@@
Algorithm 1 Decision Algorithm for Choosing Between MLP and KAN and Configuring KANs
1:Input: task type, data scale, solution structure, constraints (speed, accuracy, physics)
2:Output: model class and recommended KAN configuration
3:Step 0: Identify task and constraints
4:Determine task: regression, PDE (PINN/PIKAN), operator learning, or classification
5:Assess data scale (few vs. many samples) and structural priors
6:Specify priorities: speed vs. accuracy, interpretability, physics constraints
7:Step 1: Choose MLP or KAN
8:if data are large and no structural prior is available then
9:  Use a standard MLP baseline
10:else
11:  Select a KAN architecture
12:end if
13:Step 2: Select KAN basis by solution characteristics
14:if solution is smooth or boundary-sensitive then
15:  Use B-spline KAN (cubic, extended grids)
16:else if solution is periodic or globally oscillatory then
17:  Use FourierKAN or KAF (RFF-based)
18:else if solution is polynomial-like or spectrally smooth then
19:  Use Chebyshev or Jacobi KAN with tanh normalization
20:else if solution has shocks, discontinuities, or boundary layers then
21:  Use SincKAN, rKAN, DKAN, or HRKAN
22:else if solution is multiscale or spatially heterogeneous then
23:  Use Wavelet KAN or FBKAN (PoU decomposition)
24:else if problem is high-dimensional or separable then
25:  Use KKAN, SPIKAN, or KAF (RFF)
26:end if
27:Step 3: Configure grid and basis resolution
28:Use coarse grids initially; refine adaptively (multilevel or free knots)
29:Apply residual-adaptive sampling if residuals localize
30:Tune RBF width, wavelet scale, or Sinc step sizes as needed
31:Step 4: Add physics-informed constraints (if applicable)
32:if physics constraints or PDE residuals are present then
33:  Construct PINN/PIKAN loss with data, PDE, and BC terms
34:  Apply residual reweighting, entropy viscosity, or augmented Lagrangian
35:end if
36:Step 5: Choose optimization strategy
37:Warm up with Adam/RAdam
38:Refine with (L-)BFGS or self-scaled quasi-Newton methods
39:Use variance-preserving initialization and moderate regularization
40:Step 6: Optimize for speed or accuracy
41:if speed is the primary objective then
42:  Prefer ReLU-KAN, HRKAN, FastKAN, LeanKAN
43:  Enable JAX/CUDA kernels and domain decomposition
44:else
45:  Prefer B-spline, Chebyshev/Jacobi, Wavelet KANs
46:  Use adaptive grids, physics losses, and staged optimization
47:end if
48:Step 7: Final refinement
49:Apply pruning and sparsity regularization
50:Increase capacity locally rather than uniformly
51:Validate convergence and generalization
@@ 11 Current Gaps and Path Forward
Despite rapid growth in publications and open-source implementations (Tables LABEL:KAN_vs_MLP, LABEL:githubs), the KAN ecosystem remains fragmented. Most studies show strong empirical potential in regression, operator learning, and PDE solving, yet comparisons are inconsistent and rarely grounded in theory. The field now requires principled foundations rather than isolated “KAN vs. MLP” or “basis-vs.-basis” experiments.
@@ 12 Conclusion
This review has presented a systematic and unified perspective on Kolmogorov–Arnold Networks (KANs), bridging their theoretical foundations, architectural principles, and rapidly growing body of practical applications. By synthesizing insights from the literature together with our own analysis, we identify several key conclusions that characterize the current state of KAN research and point toward promising future directions.
Theoretical perspective.
Modern KANs are best understood as architectures inspired by, rather than direct realizations of, the Kolmogorov Superposition Theorem (KST). The KST provides a conceptual blueprint: its central insight—that any continuous multivariate function admits a representation as sums of univariate functions—motivates the characteristic activate–then–sum structure of KANs. However, contemporary KAN architectures depart fundamentally from the classical theorem by replacing non-smooth, universal inner functions with smooth, learnable basis functions such as splines or polynomials. This departure sacrifices a direct constructive link to KST, but enables efficient, stable, and data-driven approximation in practical settings.
Relationship to classical and modern methods.
By positioning KANs relative to kernel methods and multilayer perceptrons (MLPs), this review clarifies their functional role and architectural advantages. We showed that shallow, one-dimensional KANs are mathematically equivalent to classical kernel or basis-expansion methods, differing primarily in their training procedure. In contrast, higher-dimensional or deep KANs diverge structurally from kernel methods by constructing multivariate interactions through additive, compositional layers rather than explicit tensor-product or radial bases, thereby mitigating the curse of dimensionality.
The introduction of depth further distinguishes KANs from linear basis-expansion models. Hidden layers induce both a multiplicative growth in effective polynomial degree and a nonlinear coupling of the final function coefficients. This coefficient entanglement underlies the expressive power of deep KANs, but also explains the optimization challenges observed in practice, including non-convex loss landscapes, parameter redundancy, and sensitivity to initialization.
From the perspective of neural networks, KANs and MLPs are expressively equivalent, but differ markedly in how nonlinearity is deployed. KANs relocate nonlinearity from fixed node-wise activations to learnable edge-wise functions, replacing the MLP paradigm of sum–then–activate with an activate–then–sum architecture. This structural shift provides a strong inductive bias, improves parameter efficiency, and enhances interpretability through explicit functional components.
Practical design and applications.
A central lesson emerging from the expanding KAN ecosystem is that performance is not intrinsic to the “KAN” label itself, but instead arises from informed design choices. Among these, basis selection is the dominant design axis. B-splines naturally encode locality, Chebyshev polynomials offer spectral accuracy for smooth problems, Fourier bases are well suited for periodic structures, and wavelets enable multi-scale representations. No single basis is universally optimal; effective KAN design must align the chosen basis with the structure of the target function.
Beyond basis choice, successful KAN applications consistently employ a rich toolkit of accuracy and efficiency enhancements. These include physics-informed loss formulations, adaptive sampling and grid strategies, domain and function decomposition, and problem-specific regularization techniques. Such methods, often inherited from the broader scientific machine learning literature, are essential for achieving robust and competitive performance.
Outlook.
In summary, Kolmogorov–Arnold Networks represent a shift away from monolithic, fixed-activation models toward structured, modular, and interpretable architectures. Their strength lies not in a single rigid formulation, but in the flexibility to encode prior knowledge directly into the functional building blocks of the network. Future progress will depend on the development of a more unified theoretical framework that connects basis selection, optimization dynamics, conditioning, and generalization. Establishing such principles will be key to fully realizing the potential of KANs as a general-purpose tool for scientific computing and data-driven modeling.